# Not All Data are Good Labels: On the Self-supervised Labeling for Time Series Forecasting

### v2

@@ D.1 Distribution of Reconstructed Datasets in Grid Search
These results suggest that datasets with moderate reconstruction loss — neither too high nor too low — strike a better balance between flexibility and generalization. This balance ultimately leads to improved predictor performance, as overly strict reconstruction constraints may limit model adaptability, while overly high reconstruction loss may fail to capture meaningful patterns.
D.2 Demystifying the Self-supervision in Scam
@@
As mentioned in Appendix C.2, the FFN in $g(\cdot;\phi)$ is designed point-wise, decoding the concatenated outputs of the convolution layers for each point independently. To further investigate the utility of conv-concat layers, we skip the initial linear layers and activation functions in $g(\cdot;\phi)$, directly applying the final linear transformation to the latent outputs of the conv-concat layers. This process generates an intermediate series, which we term as undecided $\tilde{y}^{\prime}$.
@@
Channel-wise Distribution Alignment. Distribution shift, a core challenge in long-term Time Series Forecasting (TSF), is decisive for a TSF model to generalize on future data after training. The most widely adopted approach to address this issue is Reversible Instance Normalization (RevIn) (Kim et al., 2021; Liu et al., 2022), which aligns the distributions of historical and future data.
@@
For models designed with channel-independence (CI), such as PatchTST (Nie et al., 2022) and cycleNet (Lin et al., 2024b), aligning distributions across channels is of critical importance, especially when trained on datasets with a large number of variates. The channel-wise alignment introduced by our self-supervised reconstruction task provides an effective solution to distribution shifts, thereby enhancing the performance of CI predictor models.
@@
Appendix E Related Work Revisit
@@
Channel-independent Transformers. Time Series Transformers (TST) (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022) have recently led to significant progress in the TSF problem, demonstrating convincing superiority over traditional methods and convolution-based models. Inspired by (Zeng et al., 2023), (Nie et al., 2022) incorporates the Channel Independent (CI) design (sharing weights across all channels) to introduce PatchTST, a new state-of-the-art (SOTA) model that significantly benefits from CI and the patching operation. Multiple works follow this practice and achieve excellent performance in TSF (Das et al., 2024; Liu et al., 2024a; Zhang et al., 2024).
Channel-wise Transformers. Building on TST, more recent research has focused on designing Transformers capable of capturing channel dependencies inherent in multivariate time series data. Notable examples include Crossformer (Zhang & Yan, 2023), iTransformer (iTrans(Liu et al., 2024b), Dsformer (Yu et al., 2023), and card (Wang et al., 2024). These models shed light on exploiting inter-series relationships to improve forecasting accuracy. Furthermore, having observed that Time Series Transformers are inherently unstable due to a sharp loss scape, (Ilbert et al., 2024) propose a sharpness-aware optimizer to mitigate such issues. Their work focuses on an optimization-level approach involving a two-step backward. Nonetheless, channel-wise Transformers still suffer from overfitting on small datasets. By sharpness analysis, our proposed Scam locates the overfitting issue and provides a solution from a data-level perspective.
Linear or MLP-based Models. In contrast to the quadratic complexity of Transformers, lightweight linear or MLP-based models have emerged as competitive alternatives offering simplicity and efficiency. RLinear and RMlp (Li et al., 2023) verify that a vanilla linear model or a 2-layer MLP, when combined with a widely-adopted normalization method (Kim et al., 2021), can achieve near SOTA performance in TSF. Further research (Zhao & Shen, 2024) on channel dependencies within linear and MLP-based models has yielded performance improvements over previous CI approaches. Moreover, the models (Xu et al., 2024b; Yi et al., 2024) that directly learn linear regression or MLP-based models on complex frequency features have achieved remarkable performances. Most recently, SparseTSF (Lin et al., 2024a), a highly lightweight model, incorporates 1D convolutions as down-sampling modules and learns linear parameters on the down-sampled values of the original series. cycleNet (Lin et al., 2024b), a SOTA model that explicitly captures periodic trend features to enhance vanilla linear or MLP-based models to be on par with Transformer-based models.
Appendix F Experiment Details
F.1 Datasets
@@
  • ETT (Zhou et al., 2021): This dataset contains 7 factors of electricity transformers, recorded between July 2016 and July 2018. The subsets ETTh1 and ETTh2 are recorded hourly, while ETTm1 and ETTm2 are recorded every 15 minutes.
  • Electricity (Wu et al., 2021): This dataset records the hourly electricity consumption of 321 clients.
  • Traffic (Wu et al., 2021): This dataset collects hourly road occupancy rates measured by 862 sensors across the San Francisco Bay Area freeways, spanning from January 2015 to December 2016.
  • Weather (Wu et al., 2021): This dataset includes 21 meteorological factors, recorded every 10 minutes at the Weather Station of the Max Planck Biogeochemistry Institute in 2020.
  • PeMS: This dataset contains public traffic network data from California, collected at 5-minute intervals. We use the same four subsets (PeMS03, PeMS04, PeMS07, PeMS08) as adopted in iTransformer (iTrans(Liu et al., 2024b).
For the ETT datasets, we divide them by ratio $\{0.6,0.2,0.2\}$ into train set, validation set, and test set. For Electricity, Traffic, and Weather, we follow the same split ratio of $\{0.7,0.1,0.2\}$ as in TimesNet (Wu et al., 2023; Zhou et al., 2021; Wu et al., 2021). For the PeMS datasets, we split them using the ratio $\{0.6,0.2,0.2\}$ following the same setting as iTrans (Liu et al., 2024b). All datasets are scaled using the mean and variance of their respective training sets, a standard practice in TSF (Wu et al., 2023). The statistics of all used datasets are listed in Table 4.
@@ F.2 Backbones
    Mlp (Li et al., 2023), a 2-layer MLP model combined with Reversible Instance Normalization (Kim et al., 2021).
@@
    cycleNet (Lin et al., 2024b), a 2-layer MLP equipped with efficient cycle modeling that belongs to a general seasonal-trend decomposition method (Cycle/MLP in the original paper).
@@
    iTrans (Liu et al., 2024b), a Transformer-based model that computes attention scores on the inverted series along the channel dimension.
@@
    PatchTST (Nie et al., 2022), a Channel-Independent Transformer-based model that uses patching to tokenize the input.
@@
F.3 Reproductivity
Hyperparameters and Settings. All experiments and methods are implemented in Python and PyTorch (Paszke et al., 2019) and conducted on two Nvidia RTX A5000 Ada generation GPUs (32GB VRAMs) and two Nvidia RTX A6000 GPUs (48GB VRAMs). We use the ADAM optimizer (Kingma, 2014) with a learning rate initialized as $\eta=0.001$ for all settings. Unlike prior works (Liu et al., 2024b; Lin et al., 2024a) that set a fixed, small number of training epochs, we adopt an early stopping strategy based on the MSE metric of the validation set, with a patience of 20 epochs.
@@ F.3 Reproductivity
We use 4 convolution layers in total, with consistent settings for kernel size, stride, and padding size as explained in Appendix C.1. Dim_multiplier represents the expansion ratio of convolution channels. The channels for the four convolution layers are set to 1, 2, 4, and 8, respectively, and are further multiplied by 4 to increase the model’s capacity. Hidden_dim is the dimension size in the FFN, and # of series indicates that we use 8 point-wise linear projectors in parallel as explained in Appendix C.2 and Figure 14. The hyperparameter settings provided in Table 5 are applied consistently across all datasets.
Rerun Baselines as Additional Comparisons. To ensure fair comparisons with baseline results, we include both the results reported in the original papers and the results from our own re-implementations. Specifically: The Mlp and cycleNet results are taken from cycleNet paper (Lin et al., 2024b). The PatchTST and iTrans results are from the iTrans paper (Liu et al., 2024b), as the PatchTST paper (Nie et al., 2022) does not provide results with a look-back length of 96.
@@
The Result Discrepancy on PeMS Datasets. We also evaluate the larger traffic dataset, PeMS, which has been previously examined in both iTrans and cycleNet. However, we observe major discrepancies between our results and those reported in the original papers (Liu et al., 2024b; Lin et al., 2024b).
@@
Appendix G More Experiment Results
@@
In Table 6,  7 and  8, color RED indicates better performance and color BLUE indicates worse performance.
@@
ImpAvg. - - 3.24% 1.48% - - 0.57% 0.23% - - 19.61% 14.06% - - 2.51% 1.55%
Table 7: Full results for performance comparisons between backbone models and their integration with our proposals on ETT, Electricity, Traffic, and Weather datasets. The results for Mlp and cycleNet are taken from the cycleNet paper (Nie et al., 2022), while the results for PatchTST and iTrans are from the iTransformer paper (Liu et al., 2024b).

### v3

@@ E.1 Distribution of Reconstructed Datasets in Grid Search
These results suggest that datasets with moderate reconstruction loss — neither too high nor too low — strike a better balance between flexibility and generalization. This balance ultimately leads to improved predictor performance, as overly strict reconstruction constraints may limit model adaptability, while overly high reconstruction loss may fail to capture meaningful patterns.s
E.2 Demystifying the Self-supervision in Scam
@@
As mentioned in Appendix D.2, the FFN in $g(\cdot;\phi)$ is designed point-wise, decoding the concatenated outputs of the convolution layers for each point independently. To further investigate the utility of conv-concat layers, we skip the initial linear layers and activation functions in $g(\cdot;\phi)$, directly applying the final linear transformation to the latent outputs of the conv-concat layers. This process generates an intermediate series, which we term as undecided $\tilde{y}^{\prime}$.
Channel-wise Distribution Alignment. Distribution shift, a core challenge in long-term Time Series Forecasting (TSF), is decisive for a TSF model to generalize on future data after training. The most widely adopted approach to address this issue is Reversible Instance Normalization (RevIn) [12, 20], which aligns the distributions of historical and future data.
@@
For models designed with channel-independence (CI), such as PatchTST [27] and cycleNet [17], aligning distributions across channels is of critical importance, especially when trained on datasets with a large number of variates. The channel-wise alignment introduced by our self-supervised reconstruction task provides an effective solution to distribution shifts, thereby enhancing the performance of CI predictor models.
@@
E.3 In-depth Examination on snr
We further verify the effectiveness of snr by analyzing the loss landscape. Transformer-based models, as discussed, are more prone to overfitting, particularly on noisy datasets. This weakens our adaptive mask’s ability to discard overfitted components, as $\ell_{pred}$ — the predictor’s training loss — also contributes to overfitting. Gradient penalty offers a solution to this issue. For TSF models specifically, where both inputs and outputs are time series with abrupt distribution shifts, parameter robustness is critical for mitigating overfitting. Previous works [46, 26] have shown that regulating parameters via their spectral norm enhances stability against input perturbations, which is particularly beneficial in TSF.
Figure 16: Sharpness of difference components in iTrans.
Figure 17: iTrans as a TST example
In Figure 17, the bars represent the sharpness of different components in iTrans, divided into three parts: embedding, encoder, and projector (see Figure 17). The embedding and projector are linear layers, while the encoder comprises channel-wise attention blocks. This architectural pattern is common among Time Series Transformers (TSTs), with the encoder being the primary focus and less emphasis placed on pre- or post-encoder linear layers.
Figure 18: Empirical results on variants of snr.
Without snr, the input linear layer (embedding) exhibits the highest sharpness, indicating it is the most overfitted component (see red bars). This suggests that overfitting in Transformers originates not only from self-attention mechanisms but also from the linear layers. By applying snr to the pre- and post-encoder linear layers (Section 3.4), we observe smoother loss landscapes, validating snr’s ability to reduce sharpness effectively.
Appendix F Related Work Revisit
@@
Channel-independent Transformers. Time Series Transformers (TST) [53, 40, 54] have recently led to significant progress in the TSF problem, demonstrating convincing superiority over traditional methods and convolution-based models. Inspired by [48], [27] incorporates the Channel Independent (CI) design (sharing weights across all channels) to introduce PatchTST, a new state-of-the-art (SOTA) model that significantly benefits from CI and the patching operation. Multiple works follow this practice and achieve excellent performance in TSF [6, 19, 50].
Channel-wise Transformers. Building on TST, more recent research has focused on designing Transformers capable of capturing channel dependencies inherent in multivariate time series data. Notable examples include Crossformer [51], iTransformer (iTrans[21], Dsformer [47], and card [37]. These models shed light on exploiting inter-series relationships to improve forecasting accuracy. Furthermore, having observed that Time Series Transformers are inherently unstable due to a sharp loss scape, [10] propose a sharpness-aware optimizer to mitigate such issues. Their work focuses on an optimization-level approach involving a two-step backward. Nonetheless, channel-wise Transformers still suffer from overfitting on small datasets. By sharpness analysis, our proposed Scam locates the overfitting issue and provides a solution from a data-level perspective.
Linear or MLP-based Models. In contrast to the quadratic complexity of Transformers, lightweight linear or MLP-based models have emerged as competitive alternatives offering simplicity and efficiency. RLinear and RMlp [15] verify that a vanilla linear model or a 2-layer MLP, when combined with a widely-adopted normalization method [12], can achieve near SOTA performance in TSF. Further research [52] on channel dependencies within linear and MLP-based models has yielded performance improvements over previous CI approaches. Moreover, the models [44, 45] that directly learn linear regression or MLP-based models on complex frequency features have achieved remarkable performances. Most recently, SparseTSF [16], a highly lightweight model, incorporates 1D convolutions as down-sampling modules and learns linear parameters on the down-sampled values of the original series. cycleNet [17], a SOTA model that explicitly captures periodic trend features to enhance vanilla linear or MLP-based models to be on par with Transformer-based models.
Appendix G Experiment Details
G.1 Datasets
@@
  • ETT [53]: This dataset contains 7 factors of electricity transformers, recorded between July 2016 and July 2018. The subsets ETTh1 and ETTh2 are recorded hourly, while ETTm1 and ETTm2 are recorded every 15 minutes.
  • Electricity [40]: This dataset records the hourly electricity consumption of 321 clients.
  • Traffic [40]: This dataset collects hourly road occupancy rates measured by 862 sensors across the San Francisco Bay Area freeways, spanning from January 2015 to December 2016.
  • Weather [40]: This dataset includes 21 meteorological factors, recorded every 10 minutes at the Weather Station of the Max Planck Biogeochemistry Institute in 2020.
  • PeMS: This dataset contains public traffic network data from California, collected at 5-minute intervals. We use the same four subsets (PeMS03, PeMS04, PeMS07, PeMS08) as adopted in iTransformer (iTrans[21].
For the ETT datasets, we divide them by ratio $\{0.6,0.2,0.2\}$ into train set, validation set, and test set. For Electricity, Traffic, and Weather, we follow the same split ratio of $\{0.7,0.1,0.2\}$ as in TimesNet [41, 53, 40]. For the PeMS datasets, we split them using the ratio $\{0.6,0.2,0.2\}$ following the same setting as iTrans [21]. All datasets are scaled using the mean and variance of their respective training sets, a standard practice in TSF [41]. The statistics of all used datasets are listed in Table 4.
@@ G.2 Backbones
    Mlp [15], a 2-layer MLP model combined with Reversible Instance Normalization [12].
@@
    cycleNet [17], a 2-layer MLP equipped with efficient cycle modeling that belongs to a general seasonal-trend decomposition method (Cycle/MLP in the original paper).
@@
    iTrans [21], a Transformer-based model that computes attention scores on the inverted series along the channel dimension.
@@
    PatchTST [27], a Channel-Independent Transformer-based model that uses patching to tokenize the input.
@@
G.3 Reproductivity
Hyperparameters and Settings. All experiments and methods are implemented in Python and PyTorch [29] and conducted on two Nvidia RTX A5000 Ada generation GPUs (32GB VRAMs) and two Nvidia RTX A6000 GPUs (48GB VRAMs). We use the ADAM optimizer [13] with a learning rate initialized as $\eta=0.001$ for all settings. Unlike prior works [21, 16] that set a fixed, small number of training epochs, we adopt an early stopping strategy based on the MSE metric of the validation set, with a patience of 20 epochs.
@@ G.3 Reproductivity
We use 4 convolution layers in total, with consistent settings for kernel size, stride, and padding size as explained in Appendix D.1. Dim_multiplier represents the expansion ratio of convolution channels. The channels for the four convolution layers are set to 1, 2, 4, and 8, respectively, and are further multiplied by 4 to increase the model’s capacity. Hidden_dim is the dimension size in the FFN, and # of series indicates that we use 8 point-wise linear projectors in parallel as explained in Appendix D.2 and Figure 11. The hyperparameter settings provided in Table 5 are applied consistently across all datasets.
Rerun Baselines as Additional Comparisons. To ensure fair comparisons with baseline results, we include both the results reported in the original papers and the results from our own re-implementations. Specifically: The Mlp and cycleNet results are taken from cycleNet paper [17]. The PatchTST and iTrans results are from the iTrans paper [21], as the PatchTST paper [27] does not provide results with a look-back length of 96.
@@
The Result Discrepancy on PeMS Datasets. We also evaluate the larger traffic dataset, PeMS, which has been previously examined in both iTrans and cycleNet. However, we observe major discrepancies between our results and those reported in the original papers [21, 17].
@@
G.4 Efficiency
Table 6: The efficiency analysis of Scam. All metrics are measured on the Electricity dataset with a batch size of 32.
# of parameters Max Memory Allocated (GB) Training Time (ms/iter)
Mlp 24.8K 0.102 6.126
+Scam 52.4K 1.095 30.451
cycleNet 78.7K 0.099 6.879
+Scam 106K 1.092 30.832
PatchTST 338K 1.281 28.074
+Scam 365K 2.072 53.044
iTrans 3.3M 0.791 20.395
+Scam 3.3M 1.729 44.357
The Scam method introduces an additional lightweight reconstruction network to aid the training. The additional computation cost is only introduced in training, because the additional module is discarded like a scaffold.
For the training efficiency, the baselines we adopted are acknowledged for their low cost. We only introduce a constant increase in memory and time consumption, enabling our method to be integrated into backbones that scale in parameters.
Appendix H More Experiment Results
@@
In Table 7,  8 and  9, color RED indicates better performance and color BLUE indicates worse performance.
@@
Imp% Avg. - - 3.24% 1.48% - - 0.57% 0.23% - - 19.61% 14.06% - - 2.51% 1.55%
Table 8: Full results for performance comparisons between backbone models and their integration with our proposals on ETT, Electricity, Traffic, and Weather datasets. The results for Mlp and cycleNet are taken from the cycleNet paper [27], while the results for PatchTST and iTrans are from the iTransformer paper [21].